OpenStudy (anonymous):

Which value for x makes the sentence true? 4x = -3x + 14 x = 6 x = 2 x = 4 x = 0

OpenStudy (wolfe8):

Plug in the values for x and see.

OpenStudy (wolfe8):

Or you can solve for x

OpenStudy (the_fizicx99):

Welll combine like terms :) 4x = -3x + 14 +3x +3x 7x = 14 /7 /7 x = 2 :P